Please, Please, Please can someone help me with these questions?? I am just totally confused. This is due tommorrow and I have no idea what i am doing. I would appreciate any and all help with these.
1) A pooled estimate of the variance is a weighted average of the variance using the two sample variances and the ___________ of each variance as the weights.
2) If a baseball player's batting average is 0.340 or 34%, find the probability that the player will have a bad season and only score at most 60 hits in the 200 times at bat?
A) 12.64%
B) 11.72%
C) 50.34%
D) 13.14%
3) When the subjects are paired or matched in some way, samples are considered to be____________?
4) What y z value corresponds to the 17% of the data between the mean and the z value?
A) 1.25
B) 0.44
C) 0.52
D) 2.10
5) A sample of 400 racing cars showed that 80 cars cost over $700,000. What is the 99% condidence interval of the true proportion of cars costing over $700,000?
6) Find the probability for P(0 A) 0.4525 or 45.25%
B) 0.4554 or 45.54%
C) 0.4207 or 42.07%
D) 0.3554 or 35.54%
7) Dr. Christina Cuttleman, a nutritionist, claims that the average number of calories in a serving of popcorn is 75 with a standard deviation of 7? A sample of 50 servings of popcorn yields an average of 78 calories. Check Dr. Cuttleman's claim at a=0.05.
8) The Eagle Ridge Contractors Association claims the average price of a home in thier subdivision is $125,150 with a standard deviation of $ 7,350. A sample of 36 homes for sale in this subdivision had an average selling price of $123,550. Is there evidence that the costs of homes for sale in this subdivision are actually lower than claimed? Test a=0.05, what is the test value?
A) -1.31
B) 1.31
C) -1.52
D) 1.52
9) Joan moves into her apartment and wants to purchase a new couch. She wants to determine if there is any difference between the average cost of couches at two different stores. Test the hypothesis that there is no difference at a=0.05
Store 1 Store 2
x(sample mean) $650 $730
o(sample deviation) $61 $ 78
n 24 21

Answer by stanbon(75887)

1) A pooled estimate of the variance is a weighted average of the variance using the two sample variances and the sum of the sample sizes of each variance as the weights.
-------------------------
2) If a baseball player's batting average is 0.340 or 34%, find the probability that the player will have a bad season and only score at most 60 hits in the 200 times at bat?
Its binomial with n=200,p=0.34,0 Use binomcdf(200,0.34,60) if you have a TI calculator
Ans: 13.14%
-----------------
A) 12.64%
B) 11.72%
C) 50.34%
D) 13.14%
--------------------
3) When the subjects are paired or matched in some way, samples are considered to be dependent.
----------------------------
4) What y z value corresponds to the 17% of the data between the mean and the z value?
P(0 < z < ?) = 0.17 ; ? = InvNorm(0.67) = 0.44
A) 1.25
B) 0.44
C) 0.52
D) 2.10
-----------------------------
5) A sample of 400 racing cars showed that 80 cars cost over $700,000. What is the 99% condidence interval of the true proportion of cars costing over $700,000?
p-hat = 80/400 = 1/5 = 0.2
E = 2.58*sqrt[0.2*0.8/400] = 0.0516
99% CI: 0.2-0.0516 < p < 0.2 + 0.0516
99% CI: 0.1484 < p < 0.2516
-------------------------------
6) Find the probability for P(0)
Comment: what do you mean by P(0)?
A)= 0.4525 or 45.25%
B) 0.4554 or 45.54%
C) 0.4207 or 42.07%
D) 0.3554 or 35.54%
-----------------------------
7) Dr. Christina Cuttleman, a nutritionist, claims that the average number of calories in a serving of popcorn is 75 with a standard deviation of 7? A sample of 50 servings of popcorn yields an average of 78 calories. Check Dr. Cuttleman's claim at a=0.05.
Ho: u = 75
Ha: u is not 75
--------------------
x-bar = 78
Critical value for alpha = 5% on a two-tail T-test with df=49 is +/-1.96
Test statistic: t(78) = (78-75)/[7/sqrt(50)] = 3.0304...
----------
Conclusion: Since the test statistic is in the reject interval, Reject Ho.
The mean is not 75
---------------------------------

8) The Eagle Ridge Contractors Association claims the average price of a home in their subdivision is $125,150 with a standard deviation of $ 7,350. A sample of 36 homes for sale in this subdivision had an average selling price of $123,550. Is there evidence that the costs of homes for sale in this subdivision are actually lower than claimed? Test a=0.05, what is the test value?
-------
t(123,550) = (123550-125150)/[7350/sqrt(36)] = -1.31
A) -1.31
B) 1.31
C) -1.52
D) 1.52
--------------------------
9) Joan moves into her apartment and wants to purchase a new couch. She wants to determine if there is any difference between the average cost of couches at two different stores. Test the hypothesis that there is no difference at a=0.05
......................Store 1 Store 2
x(sample mean).........$650... $730
o(sample deviation).... $61... $78
n....................... 24.... 21
Ho: u1-u2 = 0
Ha: u1-u2 is not 0
----------------------
I ran a 2-Sample T-test with pooled samples to get:
test-statistic = -3.8563...
p-value = 0.00038...
df = 43
-------
Conclusion:Since p-value is less than alpha=5%, Reject Ho.
There is ia difference in the average cost of couches at these two stores.
===============================
Cheers,
Stan H.